OpenStudy (anonymous):

what is the equation of the line that passes through (-2,3) and is parallel to 2x+3y=6

OpenStudy (anonymous):

how do we know if a line is parallel? the slopes are the same!

OpenStudy (anonymous):

need more help from here?

OpenStudy (anonymous):

(-2,3); 2x+3y=6 2x+3y=6 -2x -2x 3y=-2x+6 3y/3=-2/3x+2 y-3=-2/3(x-(-2)) y-9=-2x/3+4 +9 +9 y=-2x/3+13/3----->this is the answer.....